OpenStudy (anonymous):

Pls integrate following Integration sign x/x^2 +1 dx

OpenStudy (anonymous):

\[\int\limits\frac x{x^2+1}dx\]let x^2+1=y 2xdx=dy xdx=dy/2 \[\frac12\int\limits\frac {dy}y=\frac12 \ln(y) +c\]\[\frac12 \ln(x^2+1)+c=\ln \sqrt{x^2+1}+c\]

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

Nopes hehehe but what if we take x common in x^2+1

OpenStudy (anonymous):

do it..then i'll understand what u want to know exactly??

OpenStudy (anonymous):

Like x(x+1/x)

OpenStudy (anonymous):

yes..thats possible

OpenStudy (anonymous):

\[\ln \sqrt{x(x+\frac1x)}+c\]